OpenStudy (jessicawade):

polynomial function help:1 question will medal

OpenStudy (jessicawade):

write a polynomial function in standard form with zeros at -3,-1 and 1

OpenStudy (campbell_st):

well if x = -3 is a zero then (x + 3) is a factor x = -1 is a zero then (x + 1) is a factor x = 1 is a zero then (x -1) is a factor so P(x) = (x +3)(x+1)(x-1) just distribute